OpenStudy (anonymous):

@Abhisar Please help me solve this question... The efficiency of a car engine is 25%.The total resistance to motion is 1000N when the car is moving at a constant velocity.If the energy in one litre of petrol is 40MJ,how far can the car travel with 1.0 litre of petrol?

OpenStudy (abhisar):

Given energy in 1 litre petrol is 40Mj and the efficiency is 25% so the car will use 10Mj of energy

OpenStudy (abhisar):

Total work done will be equal to 10Mj, Work=Force\(\times\)Distance

OpenStudy (abhisar):

set up the equation and solve it for s

OpenStudy (anonymous):

Why the car will use 10MJ of energy?

OpenStudy (abhisar):

Because efficiency is 25%

OpenStudy (anonymous):

I'm still don't understand...

OpenStudy (abhisar):

25% of 40Mj=10Mj

OpenStudy (abhisar):

The car is able to use only 25% of total fuel energy

OpenStudy (anonymous):

O...got it...I go to try 1st...

OpenStudy (abhisar):

ok

OpenStudy (anonymous):

How can i get distance?

OpenStudy (abhisar):

Force\(\times\)Distance=Work, Here work is 10Mj and Force=1000N

OpenStudy (anonymous):

That 10MJ is work?

OpenStudy (abhisar):

yes, Energy=Work

OpenStudy (anonymous):

And what is M?

OpenStudy (abhisar):

Mj=Mega joules

OpenStudy (abhisar):

1 000 000 joules

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

1000000/1000 is 1000,but the book answer is 10km...

OpenStudy (abhisar):

10Mj= 10\(\times\) 10\(^6\) = 10\(^7\) joules, now distance will be equal to 10Km

OpenStudy (abhisar):

1Mj=10\(^6\) j

OpenStudy (anonymous):

Ya...I forgot to time 10...OK...final answer is 10km...

OpenStudy (anonymous):

If just give power,Force and speed,how can i calculate the efficiency?

OpenStudy (abhisar):

I am not getting ur question ! Efficiency = Output/Input * 100

OpenStudy (anonymous):

and g

OpenStudy (anonymous):

this is the question... The power of an electrical motor is 200W.It is used to raise a load of 100N at a constant speed 0.6m/s.What is the efficiency of the motor?(Take g=10m/s2)

OpenStudy (abhisar):

Tell me the formula of power ?

OpenStudy (anonymous):

Power=FV

OpenStudy (abhisar):

Power = Work done/Time

OpenStudy (abhisar):

now 200 W means that the motor consumes a energy of 200 joules per second but pulls only 100 N per second.

OpenStudy (abhisar):

Efficiency= 100/200*100=50%

OpenStudy (anonymous):

Book Answer is 30%

OpenStudy (anonymous):

And why i cannot use P=FV?

OpenStudy (abhisar):

Yes, you will have to use P=F.V

OpenStudy (abhisar):

P=100*0.6=60W

OpenStudy (abhisar):

Now, output power will be 60W and input power is given 200W

OpenStudy (abhisar):

Efficiency will be 60/200*100=30%

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

Ok..I got it...But how can identify which is input power and output power?

OpenStudy (abhisar):

The power which is used to perform work or is supplied is input.

OpenStudy (anonymous):

OO...What is the difference between P=work done/time and P=F.S?

OpenStudy (abhisar):

Both are same ! Just two diff ways to obtain same thing.

OpenStudy (anonymous):

ok...got it...

OpenStudy (abhisar):

\(\huge\sf\color{green}{\text{✌゚\(\ddot\smile\) ✌゚}}\)

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (anonymous):

What is resistive force?

OpenStudy (abhisar):

Force which resist motion.

OpenStudy (anonymous):

If the momentum is conserved,what formula i want to use?

OpenStudy (abhisar):

Post it as new question !